Calculate the heat capacity of Hf at 200 K, given that its Debye temperature is 

242 K.


1
Expert's answer
2021-02-22T10:28:03-0500
"C_V=\frac{12\pi^4}{5}R\left(\frac{T}{T_D}\right)^3\\C_V=\frac{12\pi^4}{5}8.31\left(\frac{200}{242}\right)^3=1097\frac{J}{molK}"